# Can Transformers Learn Full Bayesian Inference in Context?

### v1

@@ Methods.
Apart from a comparison with a gold standard, we compare our ICL approach to a Laplace approximation (Daxberger et al., 2021) and different established VI methods based on automatic differentiation VI (Kucukelbir et al., 2017). For the variational distribution, we use a normal distribution with 1) a diagonal and 2) a full covariance matrix, as well as 3) a structured normal distribution with linear dependencies between the latent variables, and 4) an approach based on inverse auto-regressive flows (IAF; Kingma et al., 2016). Appendix D contains a discussion about the hyperparameters of all considered methods.
@@
For every synthetic and real-world dataset, 1000 posterior samples from each method are compared against samples from the analytical solution, if available, or from a Hamiltonian Monte Carlo (HMC) sampler with a NUTS kernel (Hoffman et al., 2014) as the gold standard. For unimodal problems, we run a single chain. For posteriors with multiple modes, we use three times the number of modes as the number of chains to capture multimodality.
Evaluation metrics
Three metrics are employed to compare samples from different approximations of the posterior distribution. The first metric is a classifier 2-sample test (C2ST; Lueckmann et al., 2021; Lopez-Paz & Oquab, 2016), where the ROC-AUC score of a random forest classifier, trained to distinguish between samples from the gold standard and the method in question, is utilized. For random forest, we use default hyperparameters, as defined in Scikit-learn (Pedregosa et al., 2011) and 10-fold cross-validation. The second metric is the maximum mean discrepancy (MMD) between the two distributions (gold-standard and each tested method) with an exponential kernel (Gretton et al., 2012). The third metric is the empirical Wasserstein-2 distance ($\mathcal{W}_{2}$; Givens & Shortt, 1984) of the two distributions, using a quadratic solver implemented in the POT library (Flamary et al., 2021).
@@ 4.1 Generalized linear models
Across seven different variants of GLMs, we find that ICL yields samples that have overall the highest agreement with the gold standard (see Table 1). Specifically on the synthetic datasets, the C2ST, MMD and $\mathcal{W}_{2}$ metrics indicate that the posterior distribution can be approximated more accurately with ICL than via variational inference. Particularly in cases where the posterior has a shape deviating from a normal distribution, ICL and HMC agree more closely than VI. For instance, in the case where a gamma prior, i.e. a skewed distribution, is used on the coefficients of a regression model, we find that ICL substantially outperforms VI both on synthetic and real-world data (see Table 2). On the real-world data, ICL still matches the performance of VI methods and has the best (or not significantly worse than the best) performance in terms of C2ST in four out of seven cases (see Appendix B.1).
Table 1: Summarized results for GLMs. Average performance of VI methods and our ICL approach on 50 synthetic and 17 real-world datasets across 7 different GLM scenarios. Comparison to the analytical solution when available and HMC otherwise. The best average result is marked in bold.
Model Synthetic Evaluation Real-World Evaluation
Laplace Approximation 1.000 2.770 2.049 1.000 2.091 0.849
VI: DiagonalNormal 0.869 1.586 1.742 0.819 0.583 0.529
VI: MultivariateNormal 0.714 1.016 1.601 0.668 0.116 0.374
VI: Structured Normal 0.711 0.929 1.580 0.664 0.109 0.370
VI: IAF 0.784 1.648 2.349 0.732 0.516 0.680
ICL 0.657 0.183 0.556 0.648 0.090 0.387
Model C2ST ($\downarrow$) MMD ($\downarrow$) $\mathcal{W}_{2}$ ($\downarrow$)
Laplace Approximation 1.000 ($\pm$ 0.000) 1.982 (±($\pm$( ± 0.126) 0.623 ($\pm$ 0.084)
VI: DiagonalNormal 0.810 ($\pm$ 0.036) 0.441 (±($\pm$( ± 0.252) 0.384 ($\pm$ 0.089)
VI: MultivariateNormal 0.711 ($\pm$ 0.038) 0.148 (±($\pm$( ± 0.093) 0.279 ($\pm$ 0.056)
VI: Structured Normal 0.705 ($\pm$ 0.032) 0.140 (±($\pm$( ± 0.081) 0.269 ($\pm$ 0.045)
VI: IAF 0.777 ($\pm$ 0.106) 0.684 (±($\pm$( ± 0.939) 0.625 ($\pm$ 0.525)
ICL 0.610 ($\pm$ 0.045) 0.046 (±($\pm$( ± 0.020) 0.242 ($\pm$ 0.038)
@@ 4.2 Factor analysis
On the factor analysis tasks, ICL has notably lower dissimilarity scores compared to the gold standard than all other considered methods in the synthetic evaluation (Table 3). Notably, an average C2ST score of 0.568 is remarkably close to the theoretical lower bound of 0.5. Regarding the real world datasets, C2ST and MMD indicate that our ICL approach yields samples most similar to the reference, while the average $\mathcal{W}_{2}$ score is substantially higher. We hypothesize that this discrepancy in the metrics might be caused by numerical issues when computing the empirical $\mathcal{W}_{2}$ distance. Furthermore, the relatively high number of latent variables in comparison to the limited number of data-points can yield overly flexible assumptions on the variational posterior causing the VI methods to overfit. While the ICL approach is well suited for cases with little data, the small number of data points is likely the cause for the poor performance of the VI methods on the FA tasks.
Table 3: Summarized results for FA: Average performance of VI methods and our ICL approach on 50 synthetic and 17 real-world datasets across 6 different FA scenarios. Comparison to HMC samples. The best average result is marked in bold.
Model Synthetic Evaluation Real-World Evaluation
Laplace Approximation 1.000 4.115 2.543 1.000 4.127 0.597
VI: DiagonalNormal 0.999 3.321 1.998 0.960 1.220 0.288
VI: MultivariateNormal 0.993 3.222 1.955 0.950 1.173 0.281
VI: Structured Normal 0.995 3.404 2.079 0.955 1.189 0.283
VI: IAF 0.987 3.226 1.973 0.902 0.969 0.251
ICL 0.568 0.057 0.409 0.751 0.673 0.583
@@ 4.3 Gaussian mixture models
Full Bayesian inference for GMMs is arguably much more challenging than for GLMs or FA. First, the generative process of GMMs involves discrete assignments to clusters, which poses a challenge not only for NUTS, but especially for VI methods. Second, the dimensionality of the posterior samples can be relatively large since for diagonal normal distributions, each component of the mixture has a mean and a variance parameter per dimension. Finally, the considered GMMs are not identifiable leading to multi-modal posterior distributions, which are impossible to perfectly approximate with the most commonly used VI methods based on normal approximations.
Due to this inherent difficulty of the GMM scenarios, we find the overall performances of all models to be worse than in the GLM and FA cases. In particular, the C2ST metric is almost saturated for the VI approaches and has a value of around 83 percent for ICL (Table 4). The MMD and $\mathcal{W}_{2}$ metrics also indicate that ICL yields samples with higher agreement with the reference than the other approaches on synthetic data. A plot of the marginals of the posterior shows high agreement between the posterior distributions of both HMC and ICL while VI is incapable of perfectly approximating a bimodal distribution and exhibits typical mode-seeking behavior (Figure 3). Note that also the VI approach based on inverse autoregressive flows, which in theory allows flexible modeling of a wide range of posterior shapes, fails to learn the bi-modality accurately from the limited number of 50 data points in this GMM scenario. This demonstrates the strength of our ICL approach in flexibly learning distributions agnostic of the provided sample size. On the real-world evaluation, the differences are similar, albeit slightly less pronounced. While C2ST and MMD are better for ICL than for VI, the $\mathcal{W}_{2}$ metric is not substantially different.
Table 4: Summarized Results for GMMs. Average performance of VI methods and our ICL approach on 50 synthetic and 17 real-world datasets across 4 different GMM scenarios. Comparison to HMC samples. The best average result is marked in bold.
Model Synthetic Evaluation Real-World Evaluation
Laplace Approximation 1.000 3.916 8.324 1.000 3.385 12.740
VI: DiagonalNormal 0.994 2.676 7.938 0.992 2.182 11.633
VI: MultivariateNormal 0.995 2.556 7.947 0.987 2.143 11.696
VI: Structured Normal 0.994 2.595 7.929 0.988 2.129 11.521
@@
ICL 0.825 0.706 4.348 0.881 1.051 10.691
Figure 3: Density plots for the marginals of the posterior for GMM scenario 1. Comparison to HMC samples a on a synthetic dataset. Only the marginals of the first three components of the mean and the variance are shown.
@@ 5 Discussion
This paper explores in-context learning for the purpose of full Bayesian inference in latent variable models. We propose to use conditional flow matching as a generic and flexible framework to approximate posterior distributions and an architecture that utilizes a transformer encoder for potentially complex conditioning on the data. We find that our ICL approach yields, on average, a closer approximation of the posterior than several state-of-the-art variational inference methods across different datasets and model setups. This does not only hold for synthetic data, but also real-world tabular datasets, emphasizing the flexibility of ICL and its applicability for full Bayesian inference.
@@
While our experiments indicate the effectiveness of ICL as a Bayesian inference method, it requires an extensive up-front training routine on modern GPU hardware. Despite ICL being consistently faster at inference time than the considered HMC methods, the overall computational burden to train our approach is much higher. As with many other ICL approaches, large datasets as a context can further become computationally very expensive.
@@
Despite its vast up-front computational cost, ICL has not only proven fundamentally transformative in the field of natural language processing (Brown et al., 2020; Touvron et al., 2023), but recently also appears to be very promising for tabular classification (Hollmann et al., 2022). Exploring the frontiers of ICL in terms of full Bayesian inference, starting from the feasibility results of this work, might therefore yield a path into similarly fertile territories.
Even though our experiments show that ICL works well despite being trained on data that is potentially very different from real-world data, the approach will only be as flexible as the data and model structures it was trained on. As a result, ICL might fail if the model, which implies the synthetic data generation, is severely misspecified. However, this is the same limitation as when misspecifying the hypothesis space of, e.g., a deep neural network or other machine learning approaches, effectively providing the model with the wrong inductive bias.
While flexible state-of-the-art sampling-based methods, such as HMC, are an efficient and highly effective reference in terms of inference for standard and statistical methods discussed in this paper, the proposed ICL approach is fundamentally more general in nature. In particular, any probabilistic model for which a generative process is conceivable can be fitted using our ICL approach—the potential for fitting models beyond the horizon of standard Bayesian methods is therefore manifold.
Acknowledgments
VF was supported by the Branco Weiss Fellowship. DR’s research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 548823575.
@@ A.1 Generalized linear models
GLMs belong to the framework of latent variable models defined by data ${\bm{x}}$ and (latent) variables ${\bm{z}}$, where the data comprises covariates and response ${\bm{x}}\vcentcolon=({\bm{u}},y)$. The variables of interest are the coefficients ${\bm{z}}\vcentcolon=\bm{\beta}$. This yields the following generative process for a set of synthetic samples $\mathcal{D}\vcentcolon=\left\{({\bm{x}}^{(i)},{\bm{z}}^{(i)})\right\}_{i=1}^{N}$ from $P^{{\bm{x}},{\bm{z}}}$:
1
2Initialize $\mathcal{D}\leftarrow\o$;
3 for $i=1,\ldots N$ do
4       draw $\bm{\beta}_{i}\sim P^{\bm{\beta}}$ ;
5       draw $\sigma^{2}_{i}\sim P^{\bm{\sigma^{2}}}$ ;
6       for $j=1,\ldots,K$ do
7             draw $\bm{u}_{i,j}\sim P^{\bm{u}}$;
8             draw $y_{i,j}\sim p(y|g^{-1}\left(\bm{u}_{i,j}^{\top}\bm{\beta}_{i}\right),\sigma^{2}_{i})$;
9
10       end for
11      set ${\bm{x}}^{(i)}\vcentcolon=\left((\bm{u}_{i,j},y_{i,j})\right)_{j=1}^{K}$;
12       set ${\bm{z}}^{(i)}\vcentcolon=\bm{\beta}_{i}$;
13       $\mathcal{D}\leftarrow\mathcal{D}\cup\left\{({\bm{x}}^{(i)},{\bm{z}}^{(i)})\right\}$;
14
15 end for
16
17
@@ 13       $\mathcal{D}\leftarrow\mathcal{D}\cup\left\{({\bm{x}}^{(i)},{\bm{z}}^{(i)})\right\}$;
Scenario 1 $\mathcal{N}(0,1)$ $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
@@
Scenario 3 $\text{Laplace}(0,1)$ $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
@@
Scenario 5 $\text{Ga}(1,1)$ $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 6 $\mathcal{N}(0,1)$ $\text{Bin}(1,\operatorname{sigmoid}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i}))$
Scenario 7 $\mathcal{N}(0,1)$ $\text{IG}(5,2)$ $\text{Ga}(\sigma^{-2}_{i}\exp{(\bm{u}_{i,j}^{\top}\bm{\beta}_{i})},\sigma^{-2}_{i}\exp(2\bm{u}_{i,j}^{\top}\bm{\beta}_{i}))$

### v2

@@ Methods.
Apart from a comparison with a gold standard, we compare our ICL approach to a Laplace approximation (LA; Daxberger et al., 2021) and different established VI methods based on automatic differentiation VI (Kucukelbir et al., 2017). For the variational distribution, we use a normal distribution with 1) a diagonal and 2) a full covariance matrix, as well as 3) a structured normal distribution with linear dependencies between the latent variables, and 4) an approach based on inverse autoregressive flows (IAF; Kingma et al., 2016). Appendix F includes a discussion regarding the hyperparameters of all considered methods.
@@
For every synthetic and real-world dataset, 1000 posterior samples from each method are compared against samples from the analytical solution, if available, or from a Hamiltonian Monte Carlo (HMC) sampler with a NUTS kernel (Hoffman et al., 2014) as the gold standard. If posteriors are unimodal, we run a single chain. In the multimodal case, we use three times the number of modes as the number of Markov chains.
Evaluation Metrics.
Three metrics are employed to compare samples from different approximations of the posterior distribution. The first metric is a classifier 2-sample test (C2ST; Lueckmann et al., 2021; Lopez-Paz & Oquab, 2016), where the ROC-AUC score of a random forest classifier, trained to distinguish between samples from the gold standard and the method in question, is utilized. For random forest, we use default hyperparameters, as defined in Scikit-learn (Pedregosa et al., 2011) and 10-fold cross-validation. We use a random forest with the given hyperparameters as a highly performative classifier in order to detect small deviations in distributions, even though this incurs the risk that the C2ST quickly saturates at a value of one, especially in high-dimensional cases. The second metric is the maximum mean discrepancy (MMD) between the two distributions (gold-standard and each tested method) with an exponential kernel (Gretton et al., 2012). The third metric is the empirical Wasserstein-2 distance ($\mathcal{W}_{2}$; Givens & Shortt, 1984) of the two distributions, as implemented in the POT library (Flamary et al., 2021).
Model C2ST MMD $\mathcal{W}_{2}$
LA 1.000 ($\pm$ 0.000) 1.982 $(\pm$ 0.126) 0.623 ($\pm$ 0.084)
VI: Diagonal 0.810 ($\pm$ 0.036) 0.441 $(\pm$ 0.252) 0.384 ($\pm$ 0.089)
VI: Full 0.711 ($\pm$ 0.038) 0.148 $(\pm$ 0.093) 0.279 ($\pm$ 0.056)
VI: Structured 0.705 ($\pm$ 0.032) 0.140 $(\pm$ 0.081) 0.269 ($\pm$ 0.045)
VI: IAF 0.777 ($\pm$ 0.106) 0.684 $(\pm$ 0.939) 0.625 ($\pm$ 0.525)
ICL (ours) 0.610 ($\pm$ 0.045) 0.046 $(\pm$ 0.020) 0.242 ($\pm$ 0.038)
@@ 4.1 Generalized Linear Models
Across seven different variants of GLMs, we find that ICL yields samples that have overall the highest agreement with the gold standard (see Table 1). Specifically on the synthetic datasets, the C2ST, MMD and $\mathcal{W}_{2}$ metrics indicate that the posterior distribution can be approximated more accurately with ICL than via variational inference.
Particularly in cases where the posterior has a shape deviating from a normal distribution, ICL and HMC agree more closely than VI. For instance, in the case where a gamma prior, i.e. a skewed distribution, is used on the coefficients of a regression model, we find that ICL substantially outperforms VI both on synthetic and real-world data (see Table 2). On the real-world data, ICL still matches the performance of VI methods and has the best (or not significantly worse than the best) performance in terms of C2ST in four out of seven cases (see Table 2). Please refer to Appendix I for the detailed experimental results summarized in Table 1.
Table 3: Summarized results for FA. Average performance of VI methods and our ICL approach on 50 synthetic and 17 real-world datasets across 6 different FA scenarios. Comparison to HMC samples. Lower is better for all metrics. The best average result is marked in bold.
Model Synthetic Real-World
C2ST MMD $\mathcal{W}_{2}$ C2ST MMD $\mathcal{W}_{2}$
LA 1.000 4.115 2.543 1.000 4.127 0.597
VI: Diagonal 0.999 3.321 1.998 0.960 1.220 0.288
VI: Full 0.993 3.222 1.955 0.950 1.173 0.281
VI: Structured 0.995 3.404 2.079 0.955 1.189 0.283
VI: IAF 0.987 3.226 1.973 0.902 0.969 0.251
ICL (ours) 0.568 0.057 0.409 0.751 0.673 0.583
@@ 4.2 Factor analysis
On the factor analysis tasks, ICL has notably lower dissimilarity scores compared to the gold standard than all other considered methods in the synthetic evaluation (Table 3). Notably, an average C2ST score of 0.568 is remarkably close to the theoretical lower bound of 0.5. Regarding the real world datasets, C2ST and MMD indicate that our ICL approach yields samples most similar to the reference, while the average $\mathcal{W}_{2}$ score is substantially higher. We hypothesize that this discrepancy in the metrics might be caused by numerical issues when computing the empirical $\mathcal{W}_{2}$ distance. Furthermore, the relatively high number of latent variables in comparison to the limited number of data-points can yield overly flexible assumptions on the variational posterior causing the VI methods to overfit. See Appendix I for the detailed experimental results summarized in Table 3.
@@ 4.3 Gaussian Mixture Models
Full Bayesian inference for GMMs is more challenging than for GLMs or FA. First, the generative process of GMMs involves discrete assignments to clusters, which poses a challenge not only for NUTS, but especially for VI methods. Second, the dimensionality of the posterior samples can be relatively large since for diagonal normal distributions, each component of the mixture has a mean and a variance parameter per dimension. Finally, the considered GMMs are not identifiable leading to multi-modal posterior distributions, which are impossible to perfectly approximate with commonly used VI methods based on Gaussian approximations.
Due to this inherent difficulty of the GMM scenarios, we find the overall performances of all models to be worse than in the GLM and FA cases. In particular, the C2ST metric is almost saturated for the VI approaches and has a value of around 83 percent for ICL (Table 4). The MMD and $\mathcal{W}_{2}$ metrics also indicate that ICL yields samples with higher agreement with the reference than the other approaches on synthetic data. A plot of the marginals of the posterior shows high agreement between the posterior distributions of both HMC and ICL while VI is incapable of perfectly approximating a bimodal distribution and exhibits typical mode-seeking behavior (Figure 3). Note that also the VI approach based on inverse autoregressive flows, which in theory allows flexible modeling of a wide range of posterior shapes, fails to learn the bi-modality accurately from the limited number of 50 data points in this GMM scenario. This demonstrates the strength of our ICL approach in flexibly learning distributions agnostic of the provided sample size. Please refer to Appendix I for the detailed experimental results summarized in Table 4.
Table 4: Summarized Results for GMMs. Average performance of VI methods and our ICL approach on 50 synthetic and 17 real-world datasets across 4 different GMM scenarios. Comparison to HMC samples in all cases. The best average result is marked in bold.
Model Synthetic Real-World
C2ST MMD $\hskip 2.84544pt\mathcal{W}_{2}$ C2ST MMD $\hskip 5.69046pt\mathcal{W}_{2}$
LA 1.000 3.916 8.324 1.000 3.385 12.740
VI: Diagonal 0.994 2.676 7.938 0.992 2.182 11.633
VI: Full 0.995 2.556 7.947 0.987 2.143 11.696
VI: Structured 0.994 2.595 7.929 0.988 2.129 11.521
@@
ICL (ours) 0.825 0.706 4.348 0.881 1.051 10.691
Figure 3: Density plots for the marginals of the posterior for GMM scenario 1. Comparison to HMC samples on a synthetic dataset whose density is depicted as a dotted line. Only the marginals of the first two components of the mean and the variance are shown. The density of the posterior obtained via HMC is depicted as a dotted line. While the ICL method aligns with the gold-standard HMC, the VI methods have a lower level of agreement and exhibit mode-seeking behaviour.
4.4 Ablations and Further Experimental Results
In this subsection, we present various ablations concerning our ICL approach to full Bayesian inference. Due to limited space, most of the results are deferred to the appendix.
Alternatives to Flow Matching.
Appendix L contains results from an ablation study using diffusion objectives instead of flow matching, while Appendix K investigates the use of a multivariate Gaussian to parametrize $Q_{\theta}^{\bm{z}|\bm{x}}$. The empirical results from these ablations strongly indicate that flow matching is essential for achieving a close approximation of the gold-standard posterior in the scenarios we consider.
Dimensionality.
In addition, we investigate the effect of the dimensionality $K$ of the latent variable $\bm{z}\in\mathbb{R}^{K}$ for all seven different GLM scenarios. The key takeaway from our results is that for $K=20$ and $K=50$, the ICL approach performs comparably to the other methods in terms of sample similarity to HMC, but does not outperform them. Please refer to Appendix O for more details. We hypothesize that a key reason for the failure to detect meaningful differences between the methods in high dimensions is due to the curse of dimensionality affecting our metrics.
Out-of-distribution Performance.
We further investigate the robustness of our method under mild distribution shifts in Appendix 9. Our results indicate that the performance of our ICL method remains relatively stable for small distribution shifts, but increasingly degrades for a larger gap between the training and testing distribution.
Predictive Performance.
Additionally, we evaluate our ICL method and all variational approaches with respect to the predictive performance of the considered GLM setups in Appendix J. Results confirm the strong performance of variational inference methods in terms of point prediction, especially for high dimensionalities, while the ICL method is generally competitive.
Architecture.
Appendix M discusses results regarding the effect of using an MLP-based architecture. Our experimental findings confirm that the transformer-based encoder performs significantly better than an equally sized MLP encoder.
The Classifier in the C2ST Metric.
Finally, we validate the choice of a random forest classifier for the C2ST metric (Appendix Q). We find that employing a nonlinear neural network and utilizing a random forest yields an overall analogous picture in terms of the performance of all methods.
@@ 5 Discussion
This paper explores in-context learning for the purpose of full Bayesian inference in latent variable models. We propose to use conditional flow matching as a generic and flexible framework to approximate posterior distributions and an architecture that utilizes a transformer encoder for potentially complex conditioning on the data. We find that our ICL approach yields a closer approximation of the posterior than several state-of-the-art variational inference methods across different datasets and model setups. This does not only hold for synthetic, but also real-world tabular datasets.
@@
While our experiments indicate the effectiveness of ICL as a Bayesian inference method, it requires an extensive up-front training routine on modern GPU hardware. Despite ICL being consistently faster at inference time than the considered HMC methods, the overall computational burden to train our approach is much higher.
Furthermore, the goal of this work is to show that ICL can effectively learn full Bayesian inference. Our experiments therefore focus on relatively simple posterior distributions where we can compare against established methods, such as HMC. Additionally, increased dimensionality of the problems considered poses a challenge to both the ICL method and the metrics we employ. Further, as with many other ICL approaches, large datasets as a context can become computationally very expensive.
@@
Despite its vast up-front computational cost, ICL has not only proven fundamentally transformative in the field of NLP (Brown et al., 2020; Touvron et al., 2023), but has recently started to transform the field of tabular machine learning (Hollmann et al., 2022). Exploring the frontiers of ICL in terms of full Bayesian inference, starting from the feasibility results of this work, might therefore lead to similarly fertile territories.
Although ICL performs well even when trained on data that may differ from real-world distributions, its flexibility is limited by the structure of the training data. If the synthetic data is highly unrealistic, ICL may fail— much like any model with a misspecified hypothesis space that imposes an unsuitable inductive bias.
While flexible state-of-the-art sampling-based methods, such as HMC, serve as an efficient and highly effective reference in terms of inference for standard and statistical methods discussed in this paper, the proposed ICL approach is fundamentally more general in nature. In particular, any probabilistic model for which a generative process is conceivable can be fitted using our ICL approach—the potential for fitting models beyond the horizon of standard Bayesian methods is therefore manifold.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgements
We thank Beste Aydemir and Svea Reuter for their valuable support and insightful comments. VF was supported by the Branco Weiss Fellowship. DR’s research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 548823575. We also gratefully acknowledge funding provided to AR by the Munich Center for Machine Learning (MCML).
@@ A.1 Generalized Linear Models
GLMs belong to the framework of latent variable models defined by data $\bm{x}$ and (latent) variables $\bm{z}$, where the data comprises covariates and response $\bm{x}\vcentcolon=(\bm{u},y)$. The variables of interest are the coefficients $\bm{z}\vcentcolon=\bm{\beta}$. This yields the following generative process for a set of synthetic samples $\mathcal{D}\vcentcolon=\left\{(\bm{x}_{i},\bm{z}_{i})\right\}_{i=1}^{N}$ from $P^{\bm{x},\bm{z}}$:
Algorithm 1 Data Generation Algorithm
0: Number of datasets $N$, number of samples per dataset $K$, distributions $P^{\bm{\beta}},P^{\bm{\sigma^{2}}},P^{\bm{u}}$,
0: A dataset $\mathcal{D}$ of input-output pairs $\left(\bm{x}_{i},\bm{z}_{i}\right)$ for $i=1,\dots,N$.
1: Initialize $\mathcal{D}\leftarrow\varnothing$
2:for $i=1\to N$ do
3:  Draw $\bm{\beta}_{i}\sim P^{\bm{\beta}}$
4:  Draw $\sigma^{2}_{i}\sim P^{\bm{\sigma^{2}}}$
5:  for $j=1\to K$ do
6:   Draw $\bm{u}_{i,j}\sim P^{\bm{u}}$
7:   Draw $y_{i,j}\sim p\!\Bigl{(}y\;\big{|}\;g^{-1}\bigl{(}\bm{u}_{i,j}^{\top}\bm{\beta}_{i}\bigr{)},\,\sigma^{2}_{i}\Bigr{)}$
8:  end for
9:  Set $\bm{x}_{i}\coloneqq\bigl{(}\,(\bm{u}_{i,j},\,y_{i,j})\bigr{)}_{j=1}^{K}$
10:  Set $\bm{z}_{i}\coloneqq\bm{\beta}_{i}$
11:  Update $\mathcal{D}\leftarrow\mathcal{D}\cup\{\bigl{(}\bm{x}_{i},\bm{z}_{i}\bigr{)}\}$
12:end for
@@
Scenario 1 $\mathcal{N}(0,1)$ - $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
@@
Scenario 3 $\text{Laplace}(0,1)$ - $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
@@
Scenario 5 $\text{Ga}(1,1)$ - $\text{IG}(5,2)$ $\mathcal{N}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i},\sigma^{2}_{i})$
Scenario 6 $\mathcal{N}(0,1)$ - - $\text{Bin}(1,\operatorname{sigmoid}(\bm{u}_{i,j}^{\top}\bm{\beta}_{i}))$
Scenario 7 $\mathcal{N}(0,1)$ - $\text{IG}(5,2)$ $\text{Ga}(\sigma^{-2}_{i}\exp{(\bm{u}_{i,j}^{\top}\bm{\beta}_{i})},\sigma^{-2}_{i}\exp(2\bm{u}_{i,j}^{\top}\bm{\beta}_{i}))$